OpenStudy (nsh4267):

If a is a positive real number, find the following values for the given function. (Simplify your answer completely.)

OpenStudy (nsh4267):

g(x)= 14c^2, what is sgrt g(a)

OpenStudy (nsh4267):

\[g(x)=14x ^{2}, what is \sqrt{g(a)}\]

OpenStudy (campbell_st):

well it is simply \[\sqrt{g(a)} = \sqrt{14a^2} = a \sqrt{14}\]

OpenStudy (confluxepic):

That is the simplified answer. @campbell_st